OpenStudy (anonymous):

An inlet pipe can fill a tank in 15 minutes. A drain can empty the tank in 18 minutes. If the tank is empty, and both the pipe and drain are open, how long will it take before the tank overflows?

OpenStudy (lgbasallote):

i believe you use the formula \[\frac tx - \frac ty = 1\] where x is the filling pipe and y is the emptying pipe. t is the total time it takes to be fll \[\frac{t}{15} - \frac{t}{18} = 1\] \[\frac{18t}{270} - \frac{15t}{270} = 1\] \[\frac{18t - 15t}{270} = 1\] \[\frac{3t}{270} = 1\] \[3t = 270\] \[t = \frac{270}{3}\] \[t = 90 \; \text{minutes}\]

OpenStudy (lgbasallote):

did you get htat?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

what didnt you get?

OpenStudy (anonymous):

what would the answer be because 90 is not the answer

OpenStudy (anonymous):

it needs to be in hours the answer needs to be in hours

OpenStudy (lgbasallote):

how many minutes in 1 hour?

OpenStudy (anonymous):

60

OpenStudy (lgbasallote):

therefore just subtract that from 90 90 - 60 = 30 therefore 90 minutes = 1 hour and 30 minutes

OpenStudy (anonymous):

that answer is wrong webassign is not accepting it as the answer. Is there another approach to take to solve this problem?

OpenStudy (lgbasallote):

try 1.5 hours

OpenStudy (anonymous):

yea thanks it works do you mind helping me out with another problem similar to this one please

OpenStudy (lgbasallote):

well i have to go now sorry

OpenStudy (anonymous):

ok thanks anyways